1. Calculate the electric field strength 0.4m from a 0.0098C charge.
1
Expert's answer
2021-01-06T13:52:43-0500

By definition, the electric field strength of a point charge is:


"E = k\dfrac{q}{r^2}"

where "q = 0.0098C" is the charge, "r = 0.4m" is the distance from the charge, "k = 9\times 10^9N\cdot m^2/C^2" is Coulomb's constant. Thus, obtain:


"E = 9\times 10^9\dfrac{0.0098}{0.4^2} \approx 5.5\times 10^8V/m"

Answer. "5.5\times 10^8V/m".